OpenStudy (jiteshmeghwal9):

Asking this question third time. now it is a challenge question from me :- What is the difference in the sum of the squares & the difference of squares of 'n' & 'a' when \LARGE{\frac{\sqrt{a^{n-2}.3^{a+2}}}{6^n.\left( n \over 2 \right)-\left( 1-{n \over 2} \right)}=\frac{1}{4\sqrt{2^n}}} & sum of 'a' & 'n' is 12:

OpenStudy (jiteshmeghwal9):

\[\LARGE{\frac{\sqrt{a^{n-2}.3^{a+2}}}{6^n.\left( n \over 2 \right)-\left( 1-{n \over 2} \right)}=\frac{1}{4\sqrt{2^n}}}\]

OpenStudy (anonymous):

HERE IS THE PROOF

mathslover (mathslover):

k wait lemme think again :(

mathslover (mathslover):

it is : \[\large{4\sqrt{2^n}}\] right?

mathslover (mathslover):

or : \[\large{4(\sqrt{2})^n}\]

mathslover (mathslover):

@experimentX , @Callisto @rsadhvika

mathslover (mathslover):

my work is over now and m totally CONFUSED ,,, : \[\large{\frac{a^{n-2}*3^{a+2}}{(6^n-1)^2}=2^{-n-2}*\frac{n^2}{4}}\]

mathslover (mathslover):

@vishweshshrimali5

OpenStudy (experimentx):

*

mathslover (mathslover):

* why?

OpenStudy (anonymous):

@jiteshmeghwal9 could u plz tell me what is the source of this problem? (which book? or which competition?)

OpenStudy (jiteshmeghwal9):

@mukushla it is the paper of N.T.S.E which is the biggest paper of middle class in india :)